OpenStudy (anonymous):

Anyone help ? Abel currently has a balance of $10,986.50 in an account he has held for 25 years. He opened the account with an initial deposit of $5,110. What is the simple interest rate on the account? 4.0% 2.1% 8.6% 4.6%

OpenStudy (luigi0210):

Do you know the formula?

OpenStudy (anonymous):

no

OpenStudy (luigi0210):

Well aren't we screwed

OpenStudy (johnweldon1993):

I = Prt.... @Luigi0210 lol You want to solve for 'r' here

OpenStudy (johnweldon1993):

I = interest earned P = principal amount r = rate t = time (in years)

OpenStudy (luigi0210):

that's what I thought the formula was but didn't think it was right :/

OpenStudy (anonymous):

@johnweldon1993 thank you :)

OpenStudy (anonymous):

now what ??

OpenStudy (johnweldon1993):

no problem! :) and see never second guess yourself @Luigi0210

OpenStudy (johnweldon1993):

first rearrange your equation to solve for 'r' I = Prt r = ...?

OpenStudy (luigi0210):

plug in and solve :P well that's my power.. second guessing myself >.<

OpenStudy (anonymous):

we don't know r ?

OpenStudy (johnweldon1993):

No, the 'r' is the interest rate ...which is what we're solving for

OpenStudy (luigi0210):

Nope,it's like looking for x.. except x is r and there are more numbers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you do the steps

OpenStudy (luigi0210):

Well the first step that I hope we can all do is set it up :l 10,986.50=5110(25)r

OpenStudy (johnweldon1993):

Sure.....so like I said first we should rearrange this equation to solve for 'r' I = prt divide both sides by pt r = I --- pt right?

OpenStudy (johnweldon1993):

So now plug in all your numbers

OpenStudy (anonymous):

this is difficult for me nvm guys.

OpenStudy (johnweldon1993):

no it's easy you'll see the steps...if you want to continue?

OpenStudy (luigi0210):

Naw come on don't give up.. you got this!! >.<

OpenStudy (johnweldon1993):

and @Luigi0210 I does not equal 10,986.50 I is the interest earned....so it would be Current balance - Principal balance

OpenStudy (anonymous):

ok so 10,986.50 - 5110??

OpenStudy (luigi0210):

Well this why I'm the support instead of the solver :l

OpenStudy (johnweldon1993):

so @kenbaby we have \[r = \frac{ I }{ pt }\] we know 'p' is (principal balance.... 5110) t is the time (25 years) and I is the interest earned.....which is the current balance - the initial balance (10,986.50 - 5110) (5876.50) so you have all your numbers to plug in \[r = \frac{ 5876.50 }{ 5110 25 }\]

OpenStudy (johnweldon1993):

Not sure if that came out (didn't load for me) so I'll write it again just in case r = 5876.50 ----------- (5110 * 25) what does 'r' =...?

OpenStudy (anonymous):

2.1??

OpenStudy (anonymous):

oops no 21.73 is what i got

OpenStudy (johnweldon1993):

not quite.....first ...what is 5110 TIMES 25?

OpenStudy (anonymous):

127750

OpenStudy (johnweldon1993):

right....and now what is 5876.50 DIVIDED by 127750?

OpenStudy (anonymous):

0.046

OpenStudy (johnweldon1993):

perfect...and now to change that into a percent...you move the decimal place over 2 places to the right....to get ...?

OpenStudy (anonymous):

4.6?

OpenStudy (johnweldon1993):

There you go :) Your answer would be 4.6% :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (johnweldon1993):

anytime! :) and don't be so quick to give up on yourself :)

OpenStudy (anonymous):

:) can you help with a few more that I am posting I am working on a study guide and have to be done by 4:00!!

OpenStudy (johnweldon1993):

I actually have to go...(already a bit late) but I'm sure anyone else would be happy to help you! :)